he aurar Based an istncal data, humberot damage hat occur ot a pahauor Sacm, per Nea is 4.Based on this vearty causirg hail steins prdradblibal, ugi rand caon ansiar t oP (thne will be mane vin a da mag - cousig Hailstarms at hi arm, duning tu naxt 3 years o)Pmon thm haut ot the next 4 years will be inee ot any damaye Causiag hailshorms)
Answer #1

Ans:

a)

Use Poisson distribution:

P(x=k)=e^{-\lambda }*\dfrac{\lambda ^k}{k!}

On average,there will be 3*0.54=1.62 storms in 3 years

So,

P(x>2)=1-P(x<=2)

=1-P(x=0)-P(x=1)-P(x=2)

=1-e-1.62*(1.620/0!+1.621/1!+1.622/2!)

=1-e-1.62*(1+1.62+1.3122)

=0.222

b)t=4/2=2

P(there is no storm within 2 years i.e. storms comes after 2 years)=P(T>2 yrs)=e-0.54*2=0.340